Russian alphabet. ISO-8859-5 character set. Small letters: а б в г д е ё ж з и й к л м н о п р с т у ф х ц ч ш щ ъ ы ь э ю я Capital letters: А Б В Г Д Е Ё Ж З И Й К Л М Н О П Р С Т У Ф Х Ц Ч Ш Щ Ъ

Ы Ь Э Ю Я German letters: abcd äüöß ABCD ÄÖÜ